# Beyond Maximum Likelihood: Variational Inequality Estimation for Generalized Linear Models

### v1

@@ Abstract
Generalized linear models (GLMs) are fundamental tools for statistical modeling, with maximum likelihood estimation (MLE) serving as the classical method for parameter inference. While MLE performs well in canonical GLMs, it can become computationally inefficient near the true parameter value. In more general settings with non-canonical or fully general link functions, the resulting optimization landscape is often non-convex, non-smooth, and numerically unstable. To address these challenges, we investigate an alternative estimator based on solving the variational inequality (VI) formulation of the GLM likelihood equations, originally proposed by Juditsky and Nemirovski [juditsky2019signal] as an alternative for solving nonlinear least-squares problems. Unlike their focus on algorithmic convergence in monotone settings, we analyze the VI approach from a statistical perspective, comparing it systematically with the MLE. We also extend the theory of VI estimators to a broader class of link functions, including non-monotone cases satisfying a strong Minty condition, and show that it admits weaker smoothness requirements than MLE, enabling faster, more stable, and less locally trapped optimization. Theoretically, we establish both non-asymptotic estimation error bounds and asymptotic normality for the VI estimator, and further provide convergence guarantees for fixed-point and stochastic approximation algorithms. Numerical experiments show that the VI framework preserves the statistical efficiency of MLE while substantially extending its applicability to more challenging GLM settings.
@@ 1 Introduction
Generalized linear models (GLMs) are a cornerstone of statistics and machine learning, offering a unified framework that extends linear regression to accommodate diverse response types through the exponential-family formulation [nelder1972generalized, mccullagh1989generalized]. By linking the conditional mean of the response to a linear predictor via a specified link function, GLMs combine interpretability and flexibility, encompassing classical models such as linear, logistic, and Poisson regression as special cases. Their broad applicability has made GLMs indispensable across scientific, engineering, and data-driven decision-making domains.
@@
The central inferential task in GLMs is to estimate the regression parameter characterizing the relationship between covariates and responses. The maximum likelihood estimator (MLE) has long served as the canonical approach, offering strong asymptotic guarantees such as efficiency and consistency under correct model specification. However, the MLE encounters limitations in modern modeling scenarios. When the link function departs from the canonical form or incorporates non-smooth or nonlinear structures designed for robustness and flexibility, the associated likelihood-based loss function can become highly ill-conditioned, non-convex, or non-differentiable. These issues often result in slow convergence, sensitivity to initialization, and numerical instability—challenges that are amplified in high-dimensional data settings. These difficulties motivate the exploration of alternative formulations that retain the interpretability of GLMs while improving computational efficiency and providing stronger theoretical guarantees.

### v2

# arXiv:2511.03087v2[stat.ME] 01 Mar 2026

@@ Abstract
Generalized linear models (GLMs) are fundamental tools for statistical modeling, with maximum likelihood estimation (MLE) serving as the classical approach for parameter inference. While MLE performs well for canonical GLMs, it can become computationally challenging in more general settings with non-canonical, non-smooth, or nonlinear link functions, where the resulting optimization landscape may be ill-conditioned, non-convex, or non-differentiable. In this paper, we study an alternative estimation framework based on variational inequalities (VIs), which formulates GLM estimation through an operator-based equilibrium condition rather than likelihood minimization. We analyze the VI estimator from a statistical perspective and establish finite-sample error bounds and asymptotic normality under mild regularity conditions, together with convergence guarantees for fixed-point and stochastic approximation algorithms. The framework accommodates a broad class of link functions, including non-canonical and non-monotone cases satisfying a strong Minty-type condition, and extends naturally to generalized additive models via basis expansion. Numerical experiments demonstrate that the VI approach achieves competitive finite-sample accuracy and improved numerical stability relative to MLE, particularly in GLMs and GAMs with non-canonical or non-smooth link functions.
@@ 1 Introduction
Generalized linear models (GLMs) are a cornerstone of statistics and machine learning, providing a unified framework that extends linear regression to accommodate diverse response types through the exponential-family formulation [17, 16]. By linking the conditional mean of the response to a linear predictor via a specified link function, GLMs combine interpretability and flexibility, encompassing classical models such as linear, logistic, and Poisson regression as special cases. Owing to this balance, GLMs have been widely used across scientific, engineering, and data-driven decision-making domains.
@@
The central inferential task in GLMs is to estimate the regression parameter characterizing the relationship between covariates and responses. Maximum likelihood estimation (MLE) has long served as the canonical approach, offering strong asymptotic guarantees such as consistency and efficiency under correct model specification. However, in modern modeling scenarios, MLE can face substantial computational challenges. When the link function departs from the canonical form or incorporates non-smooth or nonlinear structures designed for robustness and flexibility, the resulting likelihood-based objective may become ill-conditioned, non-convex, or non-differentiable. These issues often lead to slow convergence, sensitivity to initialization, and numerical instability, particularly in high-dimensional settings. Such difficulties motivate alternative formulations that retain the interpretability of GLMs while improving computational behavior and theoretical robustness.